# Gravitational Atoms from Topological Stars

### v1

@@ II.3 The hydrogenic regime $\mu r_{b}\ll 1$
which is formally the Schrödinger equation for the hydrogen atom. On the one hand, normal modes have $\frac{\omega^{2}}{\mu^{2}}-\frac{k^{2}}{\mu^{2}}-1<0$, which in (39) corresponds to having a negative binding energy. On the other hand, for such a mode to be normalizable the Coulomb potential in (39) needs to be attractive, $\frac{\omega^{2}}{\mu^{2}}-\frac{r_{b}}{r_{s}}\frac{k^{2}}{\mu^{2}}>0$. Combining these two conditions, we learn that for a KK bound state to exist one needs
@@
Below, we will show that this is actually a necessary condition in the entire parameter space, and not only when $\mu r_{b}\ll 1$. Modes with KK momentum $p\geq p_{\star}$ are necessarily radiated to infinity, so $p_{\star}$ can be understood as the KK ionization threshold. Just as in the hydrogen atom, imposing that solutions to (39) are regular as $z\to 0$ and bound at infinity sets a quantization condition in $\omega$ which, at leading order in $\mu r_{s}$, gives:
@@
with $n_{\text{H}}=1,2,3$... and $n_{\text{H}}>\ell$. This shows that normal modes of TSs have a regime that is precisely hydrogenic, thus allowing a natural interpretation of the system as a gravitational atom. This behavior is also present in the quasi-bound states of superradiant massive fields around rotating BHs Detweiler (1980b); Baumann et al. (2019a), although in that case the states are only metastable, while here they are truly stationary. Indeed, the only decaying mechanism these states have is through monochromatic electromagnetic and gravitational wave emission.
@@ III The normal mode spectrum of Topological Stars
are the ones governing the inner structure of the potential. Furthermore, from here onward, the focus is entirely on normal modes of the TS, which have a purely real frequency, $\tilde{\omega}$, Section (II.1), implying that $\epsilon$, (43), is also purely real. With this, the radial equation (15) reads
@@
$V_{eff}$ $=\frac{\tilde{k}^{2}\,(\eta-1)}{\eta\,\tilde{r}\,(\tilde{r}-1)}+\frac{(\tilde{r}\,\eta-1)\,\Lambda}{\eta\,\tilde{r}^{3}}-\frac{\sigma}{\eta\,\tilde{r}}$. (44)
@@
The next sections are devoted to studying the space of normal mode solutions to (III). These have a purely real frequency, $\tilde{\omega}$, as shown in Section (II.1). Besides the harmonic, KK and monopole numbers $\ell,p,N$, normal frequencies are also labeled by the overtone number $n$, although we will avoid writing those labels explicitly for simplicity.
@@ III.1 Existence and Classification
As stated after (43), for the remainder of this work, $\epsilon$ will be considered purely real for normal modes of the TS. Thus, translating equation (20) to the notation here, we see that the boundary conditions for bounded modes require $\epsilon<0$, whereas for $\epsilon>0$, the solutions are QNMs, previously studied in Heidmann et al. (2023a); Bianchi et al. (2023). Hence, with the scalar equation written in a Schrödinger-like way, as in Eq. (III), we may immediately infer that normal mode solutions depend on the sign of $Q$. Regions where $Q>0$ correspond to “classically allowed” regions, where the field is allowed to oscillate. Regions with $Q<0$ correspond to “classically forbidden” regions, where bound states cannot localize. Therefore, to have normal mode solutions, the potential must allow for at least one classically allowed region outside of the star, in which case, bound modes are characterized by $V_{eff}<\epsilon<0$, as in the standard square potential problem in quantum mechanics.
With this in mind, we may proceed to verify that $\sigma>0$ is a necessary condition to have normal mode solutions. To this end, assume the opposite holds, namely $\sigma<0$. Since $\epsilon<0$ is required for the existence of normal modes, and, $\Lambda\geq 0$ by definition, together with, $2>\eta>1$ and $\tilde{r}\geq 1$, we see that $Q<0$ everywhere, so there can be no such modes. Hence, taking $\sigma>0$ with $\epsilon<0$, there is always at least one classically allowed region, where we expect normal mode solutions to exist. This existence condition is precisely (2), which was derived in the hydrogenic limit, and emerged as a condition of “attractiveness” of the gravitational interaction.
We will investigate the structure of their spectrum both analytically and numerically. The case of a perturbation with no KK momentum, $\tilde{k}=0$, is significantly easier to characterize, and, we will work out the mode classification therein in full detail by studying the potential, Q(r~)Q($\tilde${r}), in (III). The addition of KK momentum, $\tilde{k}$, carried by the scalar field will be covered afterwards, whereby most of the substantially more complicated technical work will be omitted for clarity.
@@ III.2 Zero-KK modes
Additionally, in the absence of KK momentum the existence condition $\sigma=\tilde{\mu}^{2}>0$ is always satisfied, and it is guaranteed that there will be at least one classically-allowed region where normal modes can exist. It is precisely the structure of such regions that yields a sharp classification of bound states, as we discuss next. The extrema of the potential (45) are
$\tilde{r}_{min}=\frac{3}{\eta\,(1+\alpha)},\quad\tilde{r}_{max}=\frac{3}{\eta\,(1-\alpha)},$ (46)
@@
The case $\Lambda<\Lambda_{-}$ corresponds to a low angular momentum regime (recall $\Lambda=\ell(\ell+1)-(N/2)^{2}$). In that case, given a bound mode with energy $\epsilon$ (necessarily negative), there is a unique classically-allowed region where the mode can localize, which extends from the TS center $\tilde{r}=1$ to the turning point $r_{turn}$ where $\hat{Q}(r_{turn})=0$. The case $\Lambda\geq\Lambda_{-}$ is instead a moderate or large angular momentum regime, where a bound mode with energy $\epsilon$ can see either one, or two classically allowed regions disconnected by a centrifugal barrier. Such regions may contain the star’s center, or may be disconnected from it and centered instead around $\tilde{r}_{max}$. We notice that $\tilde{r}_{max}$, whenever real, always lies outside the ISCO $\tilde{r}_{max}>\tilde{r}_{ISCO}=3/\eta$, whereas $\tilde{r}_{min}$, which always satisfies $0<\tilde{r}_{min}\leq\tilde{r}_{max}$, may even be outside the spacetime domain, $\tilde{r}_{min}<1$. After these observations, we can establish the following classification based on the classically-allowed regions that modes enjoy, and their positions relative to the ISCO:
@@
This classification is illustrated schematically in Fig 4.
@@
It can be viewed as the binding energy of the mode. $\epsilon<0$ for normal modes, hence, for them $\gamma\in\mathbb{R}$, otherwise, QNMs have imaginary $\gamma$. $\gamma\to 0$ corresponds to the ionization threshold between bound and unbound modes. Our strategy will be to first determine the regions where TS and cloud modes can possibly exist and then look at their overlap to identify the mixed ones. It turns out that the phase space split, given by (48), provides a systematic way of carrying out this task. It is important to note that $\hat{Q}$ depends on the binding energy of the mode, hence the nature of a given classically-allowed region might change depending on how “high” its energy is.
@@
Therefore, given that for $\tilde{k}=0$ we are always guaranteed at least one classically-allowed region in the star’s exterior, $\hat{Q}$ needs to have at least one positive-valued maximum (local or global) for $\tilde{r}\geq 1$. We will investigate the presence of such regions in two steps, based on (48). When no local extrema exist, (48a), one can show that the potential has a global maximum at the star surface, namely:
@@
This is always located within the ISCO, $\tilde{r}_{ISCO}=3/\eta$, and, as will become clear as we progress, is the only possible classically-allowed region where the TS family can localize.
The second case, (48b), requires us to also take into account the possible local extrema of $\hat{Q}$ given by:
$\hat{Q}(\tilde{r}=\tilde{r}_{min})=\gamma_{-}^{2}-\gamma^{2},\quad\hat{Q}(\tilde{r}=\tilde{r}_{max})=\gamma_{+}^{2}-\gamma^{2},$ (53)
@@
Note, in the range where the extrema exist, we have:
@@
$\hat{Q}_{min}$ $\leq\hat{Q}_{max},\quad\mbox{for}\quad\Lambda_{-}\leq\Lambda,$ (55)
with $\hat{Q}(\tilde{r}=\tilde{r}_{min/max})=\hat{Q}_{min/max}$ and the equalities on the left are satisfied simultaneously with those on the right. Clearly, then $\hat{Q}_{min}$ is a local minimum, whereas $\hat{Q}_{max}$ is a local maximum. Their signs $\hat{Q}_{min/max}$ change according to the binding energy of the mode as follows:
@@
$\hat{Q}_{min}$ <0 <0 >0
$\hat{Q}_{max}$ <0 >0 >0
While $\hat{Q}_{min/max}$ do not exist for $\Lambda<\Lambda_{-}$, as stated after (48), and, thus, cannot affect the structure of classically-allowed regions in that regime, the other potential well, (51), where states can possibly localize, exists for:
@@
with $\hat{\Lambda}$ defined through $\left.\gamma_{0}\right|_{\Lambda=\hat{\Lambda}}=0$, and the inequality on the second line holding for $1<\eta\leq 2$ and all $\tilde{\mu}\neq 0$. It is (III.2) that allows us to deduce that there is a range of parameters where $\hat{Q}$ possesses two distinct, classically-allowed regions with bound modes possibly having support over both of them.
As already stated, $\hat{Q}_{max}$ is always located beyond the ISCO, $\tilde{r}_{ISCO}=3/\eta$, however, the same is not true for $\hat{Q}_{min}$. When the latter is negative, it acts like a centrifugal barrier between the two possible classically-allowed regions, $\hat{Q}(\tilde{r}=1)$ and $\hat{Q}_{max}$. However, when $\hat{Q}_{min}>0$, represented by the last column of (56), we have to be more careful. Accordingly, we define:
@@
It is a simple exercise to check that $\Lambda_{+}<\hat{\Lambda}$ for all $\tilde{\mu}\neq 0$ and $1<\eta\leq 2$. We can thus conclude that in the regime:
@@
the modes energies are “high” enough that they have support over both extrema, $\hat{Q}_{min/max}$ as well as over the classically-allowed region near the star surface. In fact, they are trapped in the whole region between the perfectly reflecting boundary at the star surface and the mass barrier at infinity. In this case, deciding which family does a mode belong to, reduces, with the help of (III.2), to comparing the depths of $\hat{Q}(\tilde{r}=1)$ and $\hat{Q}_{max}$, or equivalently, $\gamma_{0}$ and $\gamma_{+}$, due to (51) and (53).
With that discussion in mind and in light of the sign distribution in (56), it is $\hat{Q}_{max}$, when representing a classically-allowed region, where cloud modes localize.
We can now look back at the definitions of the three mode families and together with (51), (53), (56) and (III.2) determine the regions where they exist. To that end, we also define:
@@
followed by the transitionary region:
@@
The splits in two parts based on the value of $\eta$ come about because $\Lambda_{+}=\Lambda_{0}$ for $\eta=\frac{3}{2}$, the transition point between TS of the first and second kind. Hence, for $1<\eta<\frac{3}{2}$, $\Lambda_{0}$ is to the right of both $\Lambda_{\pm}$, whereas, for $\frac{3}{2}<\eta\leq 2$ it sits between them, as $\Lambda_{0}>\Lambda_{-}$ as long as $\tilde{\mu}\neq 0$.
We proceed with some comments on the different regions, starting with the TS modes. The region, which always localizes near the star surface, $\tilde{r}=1$, is finite for any non-infinite and non-zero scalar mass, $\tilde{\mu}$, as well as non-extremal value of $2\geq\eta>1$. This can be seen from the requirement that $\varphi>0$ for $\gamma_{0}$ to be real, which puts a bound on $\ell$ in terms of the field’s mass. There are two exceptions to this: the extremal limit, $\eta=1$ and $\Lambda=0$ - that is $\ell=0$. The latter implies that for any non-zero scalar mass, the $\ell=0$ mode’s radial function will have at least some support in this region. The former will be investigated separately because the inner boundary condition changes. In Appendix (D) we give an approximate description of the TS modes using quasimodes Holzegel and Smulevici (2014); Keir (2016) in the regime $\tilde{\mu}\gg 1$. The resulting frequency spectrum is given by (D).
The cloud modes’ region, on the other hand, closes off only asymptotically, because:
@@
It is centered on $\tilde{r}_{max}$, which, in the eikonal limit, scales as:
$\lim\limits_{\ell\to\infty}\tilde{r}_{max}=\frac{2\,\eta\,\Lambda}{\tilde{\mu}^{2}}=\frac{2\,\eta\,\ell(\ell+1)}{\tilde{\mu}^{2}}.$ (63)
@@
As for the TS modes, we have provided an approximate construction of the cloud modes using quasimodes, Appendix (C), but this time valid for $\ell\gg 1$. It recovers (63), which also matches the geodesic analysis in Section (II.2). The frequency spectrum is given by (C). It has qualitatively the same form as the real part of the frequency [^5] In fact, we have done the computation for $0<\eta<1$, determining both the real and imaginary parts of the frequency in the eikonal limit and one can reduce the result all the way to 4D Schwarzschild. in the black string case ($0<\eta<1$) for $\ell\gg 1$. In fact, the cloud modes’ properties in the eikonal regime exactly parallel those of neutral scalar clouds around static, spherically symmetric, asymptotically flat BHs in 4D.
@@
Finally, we note that for $\eta=2$, $\hat{\Lambda}=\Lambda_{+}$, however, $\Lambda_{0}>\Lambda_{-}$ for $2\geq\eta>1$ (they are equal at $\eta=3$), indicating that the mixed region never vanishes in our range of parameters. Nevertheless, this does not necessarily imply that one will always find such a mode in the spectrum for a given $\ell,\,p,\,N,\,n$, and, indeed, we have found examples of such situations. On the other hand, when a mixed mode exists, the rough distribution of its radial function between the two classically allowed regions depends on the relative height of $\hat{Q}(\tilde{r}=1)=\gamma_{0}^{2}-\gamma^{2}$ and $\hat{Q}_{max}=\gamma_{+}^{2}-\gamma^{2}$, that is, on $\gamma_{0}^{2}-\gamma_{+}^{2}$, which vanishes for $\alpha=\frac{3-\eta}{2\,\eta}$.
@@
Figure 5: Binding energy, $\gamma$, (49), of massive, scalar field modes as a function of $\ell$, in the TS case, for $\eta=1.25$, $\tilde{\mu}=7$, $\tilde{k}=0$, $N=0$, evaluated numerically for the n=0, 1, 2, 3n=0,\,1,\,2,\,3 radial overtones in cyan disks, orange diamonds, pink pentagons and green six-pointed stars, respectively. Log plot for clarity. Filled markers are pure TS modes. Empty markers correspond to cloud modes, localized outside of the timelike geodesics ISCO. Half-filled markers indicate mixed modes. The respective regions, where our analysis specifies they can exist are shaded differently for clarity. The dashed green line for the eikonal approximation, (C), and the dashed-dotted magenta line for the large scalar mass approximation, (D), are for $n=0$ only. In black, $\gamma_{g}=\sqrt{\tilde{\mu}^{2}\,\big(1-\mathcal{E}^{2}\big)}$ with $\mathcal{E}$ from (36), represents the geodesic approximation.
Figure 6: Binding energy, $\gamma$, (49), of massive, scalar field modes as a function of $\ell$, in the TS case, for $\eta=1.75$, $\tilde{\mu}=7$, $\tilde{k}=0$, $N=0$, evaluated numerically for the n=0, 1, 2, 3n=0,\,1,\,2,\,3 radial overtones in cyan disks, orange diamonds, pink pentagons and green six-pointed stars, respectively. Log plot for clarity. Filled markers are pure TS modes. Empty markers correspond to cloud modes, localized outside of the timelike geodesics ISCO. Half-filled markers indicate mixed modes. The respective regions, where our analysis specifies they can exist are shaded differently for clarity. The dashed green line for the eikonal approximation, (C), and the dashed-dotted magenta line for the large scalar mass approximation, (D), are for $n=0$ only. In black, $\gamma_{g}=\sqrt{\tilde{\mu}^{2}\,\big(1-\mathcal{E}^{2}\big)}$ with $\mathcal{E}$ from (36), represents the geodesic approximation. Note the change in order between $\Lambda_{0}$ and $\Lambda_{+}$ with respect to Fig. (5), as anticipated by the analysis in the main text (as well as the different range on the horizontal axis).

### v2

@@ II.3 The hydrogenic regime $\mu r_{b}\ll 1$
which is formally the Schrödinger equation for the hydrogen atom. Normal modes have $\frac{\omega^{2}}{\mu^{2}}-\frac{k^{2}}{\mu^{2}}-1<0$, which in Eq. (39) corresponds to having a negative binding energy. For such a mode to be normalizable the Coulomb potential in Eq. (39) must be attractive, $\frac{\omega^{2}}{\mu^{2}}-\frac{r_{b}}{r_{s}}\frac{k^{2}}{\mu^{2}}>0$. Combining these two conditions, we learn that for a KK bound state to exist one needs
@@
Below, we will show that this is actually a necessary condition in the entire parameter space, and not only when $\mu r_{b}\ll 1$. Modes with KK momentum $p\geq p_{\star}$ are necessarily radiated to infinity, so $p_{\star}$ can be understood as the KK ionization threshold. Just as in the hydrogen atom, imposing that solutions of Eq. (39) are regular as $z\to 0$ and bound at infinity sets a quantization condition in $\omega$ which, at leading order in $\mu r_{s}$, gives
@@
with $n_{\text{H}}=1,2,3$... and $n_{\text{H}}>\ell$. This shows that normal modes of TSs have a regime that is precisely hydrogenic, thus allowing a natural interpretation of the system as a gravitational atom. This behavior is also present in the quasi-bound states of superradiant massive fields around rotating BHs Detweiler (1980b); Baumann et al. (2019a), although in that case the states are only metastable, while here they are truly stationary. Indeed, the only decaying mechanism these states have is through electromagnetic and gravitational wave emission.
@@ III The normal mode spectrum of Topological Stars
are the ones governing the inner structure of the potential. Furthermore, from here onward, the focus is entirely on normal modes of the TS, which have a purely real frequency $\tilde{\omega}$ (see Section II.1), implying that $\epsilon$, defined in Eq. (43), is also purely real. With this, the radial equation (15) reads
@@
$V_{\rm eff}$ $=\frac{\tilde{k}^{2}\,(\eta-1)}{\eta\,\tilde{r}\,(\tilde{r}-1)}+\frac{(\tilde{r}\,\eta-1)\,\Lambda}{\eta\,\tilde{r}^{3}}-\frac{\sigma}{\eta\,\tilde{r}}$, (44)
@@
The next sections are devoted to studying the space of normal mode solutions of Eq. (III). These have a purely real frequency, $\tilde{\omega}$, as shown in Section II.1. Besides the harmonic, KK and monopole numbers $\ell,p,N$, normal frequencies are also labeled by the overtone number $n$, although we will avoid writing those labels explicitly for simplicity.
@@ III.1 Existence and classification
As stated below Eq. (43), for the remainder of this work, $\epsilon$ will be considered purely real for normal modes of the TS. Thus, translating Eq. (20) to the notation here, we see that the boundary conditions for bounded modes require $\epsilon<0$, whereas for $\epsilon>0$, the solutions are QNMs, previously studied in Heidmann et al. (2023a); Bianchi et al. (2023). Hence, with the scalar equation written in a Schrödinger-like way, as in Eq. (III), we may immediately infer that normal mode solutions depend on the sign of $Q$. Regions where $Q>0$ correspond to “classically allowed” regions, where the field is allowed to oscillate. Regions with $Q<0$ correspond to “classically forbidden” regions, where bound states cannot be localized. Therefore, to have normal mode solutions, the potential must allow for at least one classically allowed region outside of the star, in which case, bound modes are characterized by $V_{\rm eff}<\epsilon<0$, as in the standard square potential problem in quantum mechanics.
With this in mind, we may proceed to verify that $\sigma>0$ is a necessary condition to have normal mode solutions. To this end, assume the opposite holds, namely $\sigma<0$. Since $\epsilon<0$ is required for the existence of normal modes, and $\Lambda\geq 0$ by definition, together with $2>\eta>1$ and $\tilde{r}\geq 1$, we see that $Q<0$ everywhere, so there can be no such modes. Hence, taking $\sigma>0$ with $\epsilon<0$, there is always at least one classically allowed region, where we expect normal mode solutions to exist. This existence condition is precisely Eq. (2), which was derived in the hydrogenic limit, and emerged as a condition of “attractiveness” of the gravitational interaction.
We will investigate the structure of the spectrum both analytically and numerically. The case of a perturbation with no KK momentum, $\tilde{k}=0$, is significantly easier to characterize, and we will work out the mode classification in full detail by studying the potential Q(r~)Q($\tilde${r}) defined in Eq. (III). The addition of KK momentum, $\tilde{k}$, carried by the scalar field will be covered afterwards, whereby most of the substantially more complicated technical work will be omitted for brevity.
@@ III.2 Zero-KK modes
Additionally, in the absence of KK momentum the existence condition $\sigma=\tilde{\mu}^{2}>0$ is always satisfied, and it is guaranteed that there will be at least one classically allowed region where normal modes can exist. It is precisely the structure of such regions that yields a sharp classification of bound states, as we discuss next. The extrema of the potential (45) are
$\tilde{r}_{\rm min}=\frac{3}{\eta\,(1+\alpha)},\quad\tilde{r}_{\rm max}=\frac{3}{\eta\,(1-\alpha)},$ (46)
@@
The case $\Lambda<\Lambda_{-}$ corresponds to a low angular momentum regime: recall that $\Lambda=\ell(\ell+1)-(N/2)^{2}$. In this case, given a bound mode with energy $\epsilon$ (necessarily negative), there is a unique classically allowed region where the mode can be localized, which extends from the TS center $\tilde{r}=1$ to the turning point $r_{\rm turn}$ where $\hat{Q}(r_{\rm turn})=0$. The case $\Lambda\geq\Lambda_{-}$ is instead a moderate or large angular momentum regime, where a bound mode with energy $\epsilon$ can see either one, or two classically allowed regions disconnected by a centrifugal barrier. Such regions may contain the star’s center, or may be disconnected from it and centered instead around $\tilde{r}_{\rm max}$. We notice that $\tilde{r}_{\rm max}$, whenever real, always lies outside the ISCO, $\tilde{r}_{\rm max}>\tilde{r}_{\rm ISCO}=3/\eta$, whereas $\tilde{r}_{\rm min}$, which always satisfies $0<\tilde{r}_{\rm min}\leq\tilde{r}_{\rm max}$, may even be outside the spacetime domain, $\tilde{r}_{\rm min}<1$. After these observations, we can establish the following classification based on the classically allowed regions for the modes, and their positions relative to the ISCO:
@@
This classification is illustrated schematically in Fig. 4.
@@
It can be viewed as the binding energy of the mode. We have $\epsilon<0$ for normal modes, hence, for them $\gamma\in\mathbb{R}$, otherwise, QNMs have imaginary $\gamma$. The limit $\gamma\to 0$ corresponds to the ionization threshold between bound and unbound modes. Our strategy will be to first determine the regions where TS and cloud modes can possibly exist, and then look at their overlap to identify the mixed ones. It turns out that the phase space split given by the condition (48) provides a systematic way of carrying out this task. It is important to note that $\hat{Q}$ depends on the binding energy of the mode, hence the nature of a given classically allowed region might change depending on its energy.
@@
Therefore, given that for $\tilde{k}=0$ we are always guaranteed at least one classically allowed region in the star’s exterior, $\hat{Q}$ needs to have at least one positive-valued (local or global) maximum for $\tilde{r}\geq 1$. We will investigate the presence of such regions in two steps, based on Eq. (48). When no local extrema exist, as in Eq. (48a), one can show that the potential has a global maximum at the star surface, namely:
@@
This is always located within the ISCO, $\tilde{r}_{\rm ISCO}=3/\eta$, and, as will become clear as we progress, it is the only possible classically allowed region where the TS family can be localized.
The second case, corresponding to Eq. (48b), requires us to also take into account the possible local extrema of $\hat{Q}$ given by:
$\hat{Q}(\tilde{r}=\tilde{r}_{\rm min})=\gamma_{-}^{2}-\gamma^{2},\quad\hat{Q}(\tilde{r}=\tilde{r}_{\rm max})=\gamma_{+}^{2}-\gamma^{2},$ (53)
@@
In the range where the extrema exist, we have:
@@
$\hat{Q}_{\rm min}$ $\leq\hat{Q}_{\rm max},\quad\mbox{for}\quad\Lambda_{-}\leq\Lambda,$ (55)
with $\hat{Q}(\tilde{r}=\tilde{r}_{\rm min/max})=\hat{Q}_{\rm min/max}$ and the equalities on the left are satisfied simultaneously with those on the right. Clearly, then $\hat{Q}_{\rm min}$ is a local minimum, whereas $\hat{Q}_{\rm max}$ is a local maximum. Their signs $\hat{Q}_{\rm min/max}$ change according to the binding energy of the mode as follows:
@@
$\hat{Q}_{\rm min}$ <0 <0 >0
$\hat{Q}_{\rm max}$ <0 >0 >0
While $\hat{Q}_{\rm min/max}$ do not exist for $\Lambda<\Lambda_{-}$, as stated below Eq. (48), and, thus, cannot affect the structure of classically allowed regions in that regime, the other potential well, Eq. (51), where states can possibly be localized, exists for:
@@
with $\hat{\Lambda}$ defined through $\left.\gamma_{0}\right|_{\Lambda=\hat{\Lambda}}=0$, and the inequality on the second line holding for $1<\eta\leq 2$ and all $\tilde{\mu}\neq 0$. It is Eq. (III.2) that allows us to deduce that there is a range of parameters where $\hat{Q}$ possesses two distinct, classically allowed regions with bound modes possibly having support over both of them.
As already stated, $\hat{Q}_{\rm max}$ is always located beyond the ISCO, $\tilde{r}_{\rm ISCO}=3/\eta$, however, the same is not true for $\hat{Q}_{\rm min}$. When the latter is negative, it acts like a centrifugal barrier between the two possible classically allowed regions, $\hat{Q}(\tilde{r}=1)$ and $\hat{Q}_{\rm max}$. However, when $\hat{Q}_{\rm min}>0$, represented by the last column of Table 56, we have to be more careful. Accordingly, we define
@@
It is a simple exercise to check that $\Lambda_{+}<\hat{\Lambda}$ for all $\tilde{\mu}\neq 0$ and $1<\eta\leq 2$. We can thus conclude that in the regime
@@
the mode energies are “high” enough that they have support over both extrema, $\hat{Q}_{\rm min/max}$, as well as over the classically allowed region near the star’s surface. In fact, they are trapped in the whole region between the perfectly reflecting boundary at the star’s surface and the mass barrier at infinity. In this case, deciding which family a mode belongs to reduces, with the help of Eq. (III.2), to comparing the depths of $\hat{Q}(\tilde{r}=1)$ and $\hat{Q}_{\rm max}$, or equivalently, $\gamma_{0}$ and $\gamma_{+}$, due to Eqs. (51) and (53).
With that discussion in mind and in light of the sign distribution in Eq. (56), it is at $\hat{Q}_{\rm max}$, when representing a classically allowed region, that cloud modes can be localized.
We can now look back at the definitions of the three mode families and together with Eqs. (51), (53), (56) and (III.2) determine the regions where they exist. To that end, we also define:
@@
followed by the transition region:
@@
The splits in two parts based on the value of $\eta$ come about because $\Lambda_{+}=\Lambda_{0}$ for $\eta=\frac{3}{2}$, the transition point between TSs of the first and second kind. Hence, for $1<\eta<\frac{3}{2}$, $\Lambda_{0}$ is to the right of both $\Lambda_{\pm}$, whereas, for $\frac{3}{2}<\eta\leq 2$ it sits between them, as $\Lambda_{0}>\Lambda_{-}$ as long as $\tilde{\mu}\neq 0$.
We proceed with some comments on the different regions, starting with the TS modes. The region which is always localized near the star surface, $\tilde{r}=1$, is finite for any non-infinite and non-zero scalar mass, $\tilde{\mu}$, as well as non-extremal value of $2\geq\eta>1$. This can be seen from the requirement that $\varphi>0$ for $\gamma_{0}$ to be real, which puts a bound on $\ell$ in terms of the field’s mass. There are two exceptions to this: the extremal limit, $\eta=1$ and $\Lambda=0$ – that is, $\ell=0$. The latter implies that for any non-zero scalar mass, the $\ell=0$ mode’s radial function will have at least some support in this region. The former will be investigated separately because the inner boundary condition changes. In Appendix D we give an approximate description of the TS modes using quasimodes Holzegel and Smulevici (2014); Keir (2016) in the regime $\tilde{\mu}\gg 1$. The resulting frequency spectrum is given by Eq. (D).
The cloud modes’ region, on the other hand, closes off only asymptotically, because
@@
It is centered on $\tilde{r}_{\rm max}$, which, in the eikonal limit, scales as:
$\lim\limits_{\ell\to\infty}\tilde{r}_{\rm max}=\frac{2\,\eta\,\Lambda}{\tilde{\mu}^{2}}=\frac{2\,\eta\,\ell(\ell+1)}{\tilde{\mu}^{2}}.$ (63)
@@
As for the TS modes, we have provided an approximate construction of the cloud modes using quasimodes in Appendix C, but this time valid for $\ell\gg 1$. It recovers Eq. (63), which also matches the geodesic analysis in Section II.2. The frequency spectrum is given by Eq. (C). It has qualitatively the same form as the real part of the frequency [^4] In fact, we have done the computation for $0<\eta<1$, determining both the real and imaginary parts of the frequency in the eikonal limit, and one can reduce the result all the way to 4D Schwarzschild. in the black string case ($0<\eta<1$) for $\ell\gg 1$. In fact, the cloud modes’ properties in the eikonal regime exactly parallel those of neutral scalar clouds around static, spherically symmetric, asymptotically flat BHs in 4D.
@@
Finally, we note that for $\eta=2$, $\hat{\Lambda}=\Lambda_{+}$, however, $\Lambda_{0}>\Lambda_{-}$ for $2\geq\eta>1$ (they are equal at $\eta=3$), indicating that the mixed region never vanishes in our range of parameters. Nevertheless, this does not necessarily imply that one will always find such a mode in the spectrum for a given $\ell,\,p,\,N,\,n$, and, indeed, we have found examples of such situations. On the other hand, when a mixed mode exists, the rough distribution of its radial function between the two classically allowed regions depends on the relative height of $\hat{Q}(\tilde{r}=1)=\gamma_{0}^{2}-\gamma^{2}$ and $\hat{Q}_{\rm max}=\gamma_{+}^{2}-\gamma^{2}$, that is, on $\gamma_{0}^{2}-\gamma_{+}^{2}$, which vanishes for $\alpha=\frac{3-\eta}{2\,\eta}$.
@@
Figure 5: Binding energy $\gamma$ of massive scalar field modes as a function of $\ell$ for a TS, as defined in Eq. (49), and plotted on a log scale for clarity. We set $\eta=1.25$, $\tilde{\mu}=7$, $\tilde{k}=0$, $N=0$, and evaluate $\gamma$ numerically for the n=0, 1, 2, 3n=0,\,1,\,2,\,3 radial overtones (cyan disks, orange diamonds, pink pentagons and green six-pointed stars, respectively). Filled markers are pure TS modes. Empty markers correspond to cloud modes, localized outside of the timelike geodesics ISCO. Half-filled markers indicate mixed modes. The regions where different modes can exist are shaded differently for clarity. The dashed green line for the eikonal approximation, Eq. (C), and the dashed-dotted magenta line for the large scalar mass approximation, Eq. (D), are for $n=0$ only. In black we plot the geodesic approximation $\gamma_{g}=\sqrt{\tilde{\mu}^{2}\,\big(1-\mathcal{E}^{2}\big)}$, with $\mathcal{E}$ from Eq. (36).
Figure 6: Same as Fig. 5, but for $\eta=1.75$, $\tilde{\mu}=7$, $\tilde{k}=0$, $N=0$. Note the change in order between $\Lambda_{0}$ and $\Lambda_{+}$ with respect to Fig. 5, as anticipated by the analysis in the main text (as well as the different range on the horizontal axis).